# Cumulants in rectangular finite free probability and beta-deformed singular values

### v1

@@ 5.2 Proof of the Main Result III
which leads to
@@
$G(z)=cz^{2},\qquad g(z)=\mathbf{q}^{-\frac{1}{2}}cz.$
@@
$G^{\,\mathbf{q}^{-\frac{1}{2}};[n]}(z)=\frac{\mathbf{q}^{\frac{1}{2}n(n-1)}\cdot n!_{\mathbf{q}^{-\frac{1}{2}}}}{n!_{\mathbf{q}}}\,c^{n}z^{2n},\quad n\in\mathbb{Z}_{\geq 1},$
@@
As a result, (81) implies
@@
By taking the $\mathbf{q}^{-\frac{1}{2}}$-derivative $D_{\mathbf{q}^{-\frac{1}{2}}}$ to both sides of (82) and by the chain rule (80), we deduce
$[n]_{\mathbf{q}}\cdot a_{2n}=\frac{1-t}{1-\mathbf{q}}\left\{m_{2n}+\sum_{\ell=1}^{n-1}{\mathbf{q}^{n-\ell}m_{2\ell}\,a_{2n-2\ell}}\right\},\quad\textrm{for all }n\in\mathbb{Z}_{\geq 1}.$ (88)
@@ 7.1 High temperature beta-singular values
In [Xu-23], the author found a one-parameter deformation to the operation of $q$-rectangular free convolution (so there are two deformation parameters $q,\gamma$ in total) by considering a novel map that takes as input two $M$-tuples $\mathbf{a}=(a_{1}\geq\dots\geq a_{M}\geq 0)$, $\mathbf{b}=(b_{1}\geq\dots\geq b_{M}\geq 0)$, as well as parameters $\beta>0$ (the “inverse temperature”), $N\in\mathbb{Z}$, $N>M$, and outputs a generalized function on $M$-tuples $\mathbf{c}=(c_{1}\geq\dots\geq c_{M}\geq 0)$. The map was regarded by the author as a “$\beta$-generalization of addition of rectangular matrices”, because when $\beta=2$, it can be interpreted as what happens when considering singular values of sums of independent $M\times N$ random matrices. To be more precise, the map mentioned originates by regarding the function $\mathbb{C}^{N\times M}\ni X\mapsto\mathbb{E}\big{[}\exp(\Re\,Tr(AX))\big{]}$ as a Fourier-transform of $A$, since for the sum $C=A+B$ of independent matrices, one has

### v3

@@ 5.2 Proof of the Main Result III
where $G(z)=\sum_{\ell=1}^{\infty}{\frac{\kappa_{2\ell}}{[\ell]_{\mathbf{q}}}\mathbf{q}^{\ell-\frac{1}{2}}z^{2\ell}}$, which leads to
@@
$G(z)=\sum_{\ell=1}^{\infty}{\frac{\kappa_{2\ell}}{[\ell]_{\mathbf{q}}}\mathbf{q}^{\ell-\frac{1}{2}}z^{2\ell}}=cz^{2},\qquad g(z)=\sum_{\ell=1}^{\infty}{\kappa_{2\ell}z^{2\ell-1}}=\mathbf{q}^{-\frac{1}{2}}cz.$
@@
$G^{\,\mathbf{q}^{-\frac{1}{2}};[n]}(z)=\frac{\mathbf{q}^{\frac{1}{2}n(n-1)}\cdot n!_{\mathbf{q}^{-\frac{1}{2}}}}{n!_{\mathbf{q}}}\,c^{n}z^{2n},\quad\text{for all }n\in\mathbb{Z}_{\geq 1},$
@@
The left hand side above is the same as the left hand side in Eqn. (82), therefore the right hand sides must also agree; as a result, by comparing the coefficients of $z^{2n}$, we find that
@@
By taking the $\mathbf{q}^{-\frac{1}{2}}$-derivative $D_{\mathbf{q}^{-\frac{1}{2}}}$ to both sides of (83) and applying the chain rule (81), we obtain
$\left(1+\sum_{i=1}^{\infty}{a_{2i}z^{2i}}\right)\cdot\left(\frac{1-t}{1-\mathbf{q}}\sum_{k=1}^{\infty}{\frac{m_{2k}}{[k]_{\mathbf{q}}}D_{\mathbf{q}^{-\frac{1}{2}}}z^{2k}}\right)=\sum_{n=1}^{\infty}{a_{2n}\,D_{\mathbf{q}^{-\frac{1}{2}}}z^{2n}}$
By using $D_{\mathbf{q}^{-\frac{1}{2}}}z^{2k}=\mathbf{q}^{-k+\frac{1}{2}}[k]_{\mathbf{q}}\cdot(1+\mathbf{q}^{\frac{1}{2}})$ and $D_{\mathbf{q}^{-\frac{1}{2}}}z^{2n}=\mathbf{q}^{-n+\frac{1}{2}}[n]_{\mathbf{q}}\cdot(1+\mathbf{q}^{\frac{1}{2}})$, it follows that
$\frac{1-t}{1-\mathbf{q}}\left(1+\sum_{i=1}^{\infty}{a_{2i}z^{2i}}\right)\left(\sum_{k=1}^{\infty}{m_{2k}\,\mathbf{q}^{-k+\frac{1}{2}}z^{2k-1}}\right)=\sum_{n=1}^{\infty}{a_{2n}\,\mathbf{q}^{-n+\frac{1}{2}}[n]_{\mathbf{q}}\,z^{2n-1}}.$
By comparing the coefficients of $z^{2n-1}$ on both sides, we deduce
$\frac{1-t}{1-\mathbf{q}}\left\{m_{2n}+\sum_{\ell=1}^{n-1}{\mathbf{q}^{n-\ell}m_{2\ell}\,a_{2n-2\ell}}\right\}=[n]_{\mathbf{q}}\cdot a_{2n},\quad\textrm{for all }n\in\mathbb{Z}_{\geq 1}.$ (89)
@@ 7.1 High temperature beta-singular values
In [Xu-23], the author found a one-parameter deformation to the operation of $q$-rectangular free convolution (so there are two deformation parameters $q,\gamma$ in total) by considering a novel map that takes as input two $M$-tuples $\mathbf{a}=(a_{1}\geq\dots\geq a_{M}\geq 0)$, $\mathbf{b}=(b_{1}\geq\dots\geq b_{M}\geq 0)$, as well as parameters $\beta>0$ (the “inverse temperature”), $N\in\mathbb{Z}$, $N>M$, and outputs a generalized function on $M$-tuples $\mathbf{c}=(c_{1}\geq\dots\geq c_{M}\geq 0)$. The map was regarded by the author as a $\beta$-generalization of addition of rectangular matrices”, because when $\beta=2$, it can be interpreted as what happens when considering singular values of sums of independent $M\times N$ random matrices. To be more precise, the map mentioned originates by regarding the function $\mathbb{C}^{N\times M}\ni X\mapsto\mathbb{E}\big[\exp(\Re\,Tr(AX))\big]$ as a Fourier-transform of $A$, since for the sum $C=A+B$ of independent matrices, one has